Provide a structure for the following compound: C10H13NO2; IR: 3285, 1659, 1246 cm?1; 1H NMR spectrum:
Answer #1

General guidance

Concepts and reason

The concept is to identify the structure of the compound from given 1HNMR^{\\rm{1}}{\\rm{HNMR}}spectrum.

Fundamentals

Number of signals in the1HNMR^{\\rm{1}}{\\rm{HNMR}}spectrum is equal to number of non-equivalent hydrogens. If proton attached to more electronegative atom (i.e. O, N, X), then it gives signal at more downfield from the reference compound (i.e. TMS). IR spectrum is useful to identify the functional groups.

Step-by-step

Step 1 of 2

Calculate the double bond equivalence is as follows.

DBE=NC+1\u2212((NH+NCl)\u2212NN2)ThemolecularformulaforthegivencompoundisC10H13NO2DBE=10+1\u2212((13+0)\u221212)DBE=5So,thecompoundhas5doublebonds,itindicatesthepresenceofbenzenering.\\begin{array}{l}\\\\{\\rm{DBE}}\\,{\\rm{ = }}\\,{{\\rm{N}}_{\\rm{C}}}\\,{\\rm{ + }}\\,{\\rm{1}}\\, - \\left( {\\frac{{{\\rm{(}}{{\\rm{N}}_{\\rm{H}}}\\,{\\rm{ + }}\\,{{\\rm{N}}_{{\\rm{Cl}}}}{\\rm{)}} - {{\\rm{N}}_{\\rm{N}}}}}{{\\rm{2}}}} \\right)\\\\\\\\{\\rm{The}}\\,{\\rm{molecular}}\\,{\\rm{formula}}\\,{\\rm{for}}\\,{\\rm{the}}\\,{\\rm{given}}\\,{\\rm{compound}}\\,{\\rm{is}}\\,\\,{{\\rm{C}}_{10}}{{\\rm{H}}_{{\\rm{13}}}}{\\rm{N}}{{\\rm{O}}_2}\\\\\\\\{\\rm{DBE}}\\,{\\rm{ = }}\\,10{\\rm{ + 1}}\\, - \\left( {\\frac{{{\\rm{(13}}\\,{\\rm{ + }}\\,{\\rm{0)}} - 1}}{{\\rm{2}}}} \\right)\\\\\\\\{\\rm{DBE}}\\,{\\rm{ = }}\\,\\,5\\\\\\\\{\\rm{So,}}\\,{\\rm{the}}\\,{\\rm{compound}}\\,{\\rm{has}}\\,\\,5\\,{\\rm{double}}\\,{\\rm{bonds,}}\\,{\\rm{it}}\\,{\\rm{indicates}}\\,{\\rm{the}}\\,{\\rm{presence}}\\,{\\rm{of}}\\,{\\rm{benzene}}\\,{\\rm{ring}}.\\\\\\end{array}

Explanation | Hint for next step

So, the compound has 5 double bonds, it indicates the presence of benzene ring.

Step 2 of 2

Total 6 signals are observed in the given 1HNMR^{\\rm{1}}{\\rm{HNMR}}spectrum.

Two 2H doublets around 6.9 to 7.5 ppm corresponds a disubstituted benzene ring.

2H quartet at 4.0 ppm and 3H triplet at 1.3 ppm are adjacent to each other.

3H singlet at 2.1 ppm means that the methyl group attached to carbonyl group.

Thus, the structure of unknown compound is,

Thus, the structure of unknown compound is,


Explanation

The IR value 1659cm\u22121{\\rm{1659}}\\;{\\rm{c}}{{\\rm{m}}^{ - {\\rm{1}}}}shows that the compound has amide linkage (-CONH2).

The IR value 1246cm\u22121{\\rm{1246}}\\;{\\rm{c}}{{\\rm{m}}^{ - {\\rm{1}}}}corresponds C-O stretching frequency.

The IR value 3285cm\u22121{\\rm{3285}}\\;{\\rm{c}}{{\\rm{m}}^{ - {\\rm{1}}}}is N-H stretching frequency.

From 1HNMR^{\\rm{1}}{\\rm{HNMR}}spectrum, triplet at 1.1 ppm and quartet is at 4.0 ppm means ethyl group is attached to the oxygen atom. Two doublets at around 6.9 to 7.5 ppm means disubstituted benzene ring. So, from all observation, the name of compound is 1,1-dichloro-2,2-diethoxyethane.

Answer

Thus, the structure of unknown compound is,